OpenStudy (lgbasallote):

\[\begin{array}{cc} \text{Find} \; \mathcal L \lbrace B(t) \rbrace \; \text{where} B(t) &= \sin2t \quad 0 \pi \end{array}\]

OpenStudy (lgbasallote):

i got something like \[\Large \left[\frac{-e^{-st} (s \sin(2t) + 2\cos (2t) )}{s^2+4} \right|_0^\pi\] is that right?

OpenStudy (anonymous):

@lgbasallote yes you are right !

OpenStudy (lgbasallote):

YAY!!! >:)))

OpenStudy (lgbasallote):

so what do i do now o.O

OpenStudy (anonymous):

just apply the limits . sin will get zero and cos will get -1

OpenStudy (lgbasallote):

i got \[\huge \frac{-2e^{-\pi s} + 2}{s^2 + 4}\] ??

OpenStudy (anonymous):

great !!!!! correct.

OpenStudy (lgbasallote):

WOOHOO!!!! \m/